OpenStudy (anonymous):

Megan has a bag of beads. She uses 28 beads to make each necklace. She makes 12 necklaces. Megan has 135 beads left over. How many beads did Megan start with?

OpenStudy (anonymous):

Pls anyone

OpenStudy (anonymous):

pls

OpenStudy (anonymous):

;(

OpenStudy (luckycoins888):

Total beads used for 12 necklace=28×12 =336 Total beads Megan start with=336+135 =471

OpenStudy (anonymous):

:o THXXXX